# Accelerating Quantum Reinforcement Learning with a Quantum Natural Policy Gradient Based Approach

### v1

@@ Appendix D Proof of Lemma 5
Since Algorithm 2 outputs an unbiased estimate of the input, we have that for any $\theta$, $\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]=E[\tilde{g}_{h}]$ and $\mathbb{E}[\hat{F}_{\rho}(\tau_{N}|\theta)]=E[\tilde{F}_{h}]$. Thus, (31) and (33) directly follow from (26) and (28). For (32) and (34), we note that
$\mathbb{E}\big{[}||\tilde{g}_{h}-\nabla_{\theta}J_{\rho}(\theta)||^{2}\big{]}=\mathbb{E}\big{[}||\tilde{g}_{h}-\mathbb{E}(\tilde{g}_{h})+\mathbb{E}(\tilde{g}_{h})-\nabla_{\theta}J_{\rho}(\theta)||^{2}\big{]}=\mathbb{E}\big{[}||\tilde{g}_{h}-\mathbb{E}(\tilde{g}_{h})||^{2}\big{]}+||\mathbb{E}(\tilde{g}_{h})-\nabla_{\theta}J_{\rho}(\theta)||^{2}$ (55)
since the cross term contain $\mathbb{E}[\tilde{g}_{h}-\mathbb{E}(\tilde{g}_{h})]$, which equals to zero. We further have
$||\mathbb{E}(\tilde{g}_{h})-\nabla_{\theta}J_{\rho}(\theta)||^{2}\leq G^{2}\bigg{(}\frac{N+1}{1-\gamma}+\frac{\gamma}{(1-\gamma)^{2}}\bigg{)}^{2}\gamma^{2N}\;\text{and}\;\mathbb{E}\big{[}||\tilde{g}_{h}-\mathbb{E}(\tilde{g}_{h})||^{2}\big{]}\leq\hat{\sigma}_{g}^{2}$ (56)
@@
$\mathbb{E}\big{[}||\tilde{F}_{h}-F_{\rho}(\theta)||^{2}\big{]}=\mathbb{E}\big{[}||\tilde{F}_{h}-\mathbb{E}(\tilde{F}_{h})+\mathbb{E}(\tilde{F}_{h})-F_{\rho}(\theta)||^{2}\big{]}=\mathbb{E}\big{[}||\tilde{F}_{h}-\mathbb{E}(\tilde{F}_{h})||^{2}\big{]}+||\mathbb{E}(\tilde{F}_{h})-F_{\rho}(\theta)||^{2}$ (57)
since the cross term contain $\mathbb{E}[\tilde{F}_{h}-\mathbb{E}(\tilde{F}_{h})]$, which equals to zero. We further have
$||\mathbb{E}(\tilde{F}_{h})-F_{\rho}(\theta)||^{2}\leq\frac{G^{4}\gamma^{2N}}{(1-\gamma)^{2}}\;\text{and}\;\mathbb{E}\big{[}||\tilde{F}_{h}-\mathbb{E}(\tilde{F}_{h})||^{2}\big{]}\leq\hat{\sigma}_{F}^{2}$ (58)
@@ Appendix E Proof of Lemma 6
where $(a)$ and $(b)$ follow from $\mu_{F}\leq{\|F_{\rho}(\theta_{k})\|}\leq G^{2}$. Taking conditional expectation on both sides, we obtain
$\mathbb{E}\left[\left\|\omega_{h+1}-\omega_{k}^{*}\right\|^{2}\right]$ $\leq(1-2\alpha\mu_{F}+2G^{4}\alpha^{2})\|\omega_{h}-\omega_{k}^{*}\|^{2}-2\alpha\langle\omega_{h}-\omega_{k}^{*},\mathbb{E}\left[\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})\right]\rangle$
$\hskip 28.45274pt+2\alpha^{2}\mathbb{E}\left\|\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})\right\|^{2}$ (59)
@@
where the last inequality follows from ${\|\omega_{k}^{*}\|}^{2}={\|F_{\rho}(\theta_{k})^{-1}\nabla_{\theta}J_{\rho}(\theta_{k})\|}^{2}\leq\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}$. Taking expectation yields
$\mathbb{E}\|\nabla_{\omega}\tilde{L}_{h}$ $-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})\|^{2}$
$\leq 3\mathbb{E}\|\tilde{F}_{h}-F_{\rho}(\theta_{k})\|^{2}\|\omega_{h}-\omega_{k}^{*}\|^{2}+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\mathbb{E}\|\tilde{F}_{h}-F_{\rho}(\theta_{k})\|^{2}+3\mathbb{E}\|\nabla_{\theta}J_{\rho}(\theta_{k})-\tilde{g}_{h}\|^{2}$
$\leq 3\left[\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right]{\|\omega_{h}-\omega_{k}^{*}\|}^{2}+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(3\hat{\sigma}_{g}^{2}+3\hat{\delta}_{g}^{2}\right)$ (60)
@@
$-\langle\omega_{h}-\omega_{k}^{*},$ $\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})]\rangle$
$\leq\frac{\mu_{F}}{4}\|\omega_{h}-\omega_{k}^{*}\|^{2}+\frac{1}{\mu_{F}}\left\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})]\right\|^{2}$
$\leq\frac{\mu_{F}}{4}\|\omega_{h}-\omega_{k}^{*}\|^{2}+\dfrac{1}{\mu_{F}}\left\|\left\{\mathbb{E}[\tilde{F}_{h}]-F_{\rho}(\theta_{k})\right\}\omega_{h}+\bigg{\{}\nabla_{\theta}J_{\rho}(\theta_{k})-\mathbb{E}\left[\tilde{g}_{h}\right]\bigg{\}}\right\|^{2}$
@@
where the last inequality follows from ${\|\omega_{k}^{*}\|}^{2}={\|F_{\rho}(\theta_{k})^{-1}\nabla_{\theta}J_{\rho}(\theta_{k})\|}^{2}\leq\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}$. Substituting the above bounds in (E), we have
$\mathbb{E}\left[\|\omega_{h+1}-\omega_{k}^{*}\|^{2}\right]$ $\leq\left(1-\frac{3\alpha\mu_{F}}{2}+\dfrac{8\alpha\hat{\delta}_{F}^{2}}{\mu_{F}}+6\alpha^{2}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+2\alpha^{2}G^{4}\right)\|\omega_{h}-\omega_{k}^{*}\|^{2}+\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]$
@@
$\mathbb{E}[\|\omega_{h+1}-\omega_{k}^{*}\|^{2}]\leq\left(1-{\alpha\mu_{F}}\right)\|\omega_{h}-\omega_{k}^{*}\|^{2}+\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha^{2}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]$
@@
$\mathbb{E}[\|\omega_{H}-\omega_{k}^{*}\|^{2}]$
$\leq\left(1-{\alpha\mu_{F}}\right)^{H}\mathbb{E}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\sum_{h=0}^{H-1}\left(1-{\alpha\mu_{F}}\right)^{h}\left\{\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha^{2}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\right\}$
$\leq\exp\left(-{H\alpha\mu_{F}}\right)\mathbb{E}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\frac{1}{\alpha\mu_{F}}\left\{\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha^{2}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\right\}$
$=\exp\left(-{H\alpha\mu_{F}}\right)\mathbb{E}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\bigg{\{}4\mu_{F}^{-2}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha\mu_{F}^{-1}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\bigg{\}}$ (62)
@@
$\|\mathbb{E}[\omega_{h+1}]-\omega_{k}^{*}\|^{2}=\|\mathbb{E}[\omega_{h}]-\alpha\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-\omega_{k}^{*}\|^{2}$
@@
$=\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\mathbb{E}[\omega_{h}]-\omega_{k}^{*},\mathbb{E}[{g}_{h}]\rangle+\alpha^{2}\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]\|^{2}$
@@
$\overset{}{=}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\mathbb{E}[\omega_{h}]-\omega_{k}^{*},F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$-2\alpha\langle\mathbb{E}[\omega_{h}]-\omega_{k}^{*},\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\rangle+\alpha^{2}\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]\|^{2}$
@@
$\overset{(a)}{\leq}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\mu_{F}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\mathbb{E}[\omega_{h}]-\omega_{k}^{*},\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$\hskip 28.45274pt+2\alpha^{2}\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\|^{2}+2\alpha^{2}\|F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\|^{2}$
@@
$\overset{(b)}{\leq}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\mu_{F}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\mathbb{E}[\omega_{h}]-\omega_{k}^{*},\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$\hskip 28.45274pt+2\alpha^{2}\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\|^{2}+2G^{4}\alpha^{2}\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}$
@@
$\leq(1-2\alpha\mu_{F}+2G^{4}\alpha^{2})\|\mathbb{E}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\mathbb{E}[\omega_{h}]-\omega_{k}^{*},\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$\hskip 28.45274pt+2\alpha^{2}\left\|\mathbb{E}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\right\|^{2}$ (63)
@@
where $(a)$ and $(b)$ follow from $\mu_{F}\leq{\|F_{\rho}(\theta_{k})\|}\leq G^{2}$. The third term in the last line of (63) can be bounded as follows.
$\|\mathbb{E}[\nabla_{\omega}$ $\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\mathbb{E}[\omega_{h}]-\omega_{k}^{*})\|^{2}$
@@
$=\left\|\mathbb{E}\left[(\tilde{F}_{h}-F_{\rho}(\theta_{k}))(\omega_{h}-\omega_{k}^{*})\right]+(\mathbb{E}[\tilde{F}_{h}]-F_{\rho}(\theta_{k}))\omega_{k}^{*}+(\nabla_{\theta}J_{\rho}(\theta_{k})-\mathbb{E}[\tilde{g}_{h}])\right\|^{2}$
@@
$\leq 3\mathbb{E}\left[\|\mathbb{E}[\tilde{F}_{h}]-F_{\rho}(\theta_{k})\|^{2}\|\omega_{h}-\omega_{k}^{*}\|^{2}\right]+3\mathbb{E}\left[\|\mathbb{E}[\tilde{F}_{h}]-F_{\rho}(\theta_{k})\|^{2}\right]\|\omega_{k}^{*}\|^{2}+3\left\|\nabla_{\theta}J_{\rho}(\theta_{k})-\mathbb{E}[\tilde{g}_{h}]\right\|^{2}$

### v3

@@ Appendix D Proof of Lemma 5
Since Algorithm 2 outputs an unbiased estimate of the input, we have that for any $\theta$, $\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]=E[\tilde{g}_{h}]$ and $\operatorname{\mathbb{E}}[\hat{F}_{\rho}(\tau_{N}|\theta)]=E[\tilde{F}_{h}]$. Thus, (31) and (33) directly follow from (26) and (28). For (32) and (34), we note that
$\operatorname{\mathbb{E}}\big{[}||\tilde{g}_{h}-\nabla_{\theta}J_{\rho}(\theta)||^{2}\big{]}=\operatorname{\mathbb{E}}\big{[}||\tilde{g}_{h}-\operatorname{\mathbb{E}}(\tilde{g}_{h})+\operatorname{\mathbb{E}}(\tilde{g}_{h})-\nabla_{\theta}J_{\rho}(\theta)||^{2}\big{]}=\operatorname{\mathbb{E}}\big{[}||\tilde{g}_{h}-\operatorname{\mathbb{E}}(\tilde{g}_{h})||^{2}\big{]}+||\operatorname{\mathbb{E}}(\tilde{g}_{h})-\nabla_{\theta}J_{\rho}(\theta)||^{2}$ (55)
since the cross term contain $\operatorname{\mathbb{E}}[\tilde{g}_{h}-\operatorname{\mathbb{E}}(\tilde{g}_{h})]$, which equals to zero. We further have
$||\operatorname{\mathbb{E}}(\tilde{g}_{h})-\nabla_{\theta}J_{\rho}(\theta)||^{2}\leq G^{2}\bigg{(}\frac{N+1}{1-\gamma}+\frac{\gamma}{(1-\gamma)^{2}}\bigg{)}^{2}\gamma^{2N}\;\text{and}\;\operatorname{\mathbb{E}}\big{[}||\tilde{g}_{h}-\operatorname{\mathbb{E}}(\tilde{g}_{h})||^{2}\big{]}\leq\hat{\sigma}_{g}^{2}$ (56)
@@
$\operatorname{\mathbb{E}}\big{[}||\tilde{F}_{h}-F_{\rho}(\theta)||^{2}\big{]}=\operatorname{\mathbb{E}}\big{[}||\tilde{F}_{h}-\operatorname{\mathbb{E}}(\tilde{F}_{h})+\operatorname{\mathbb{E}}(\tilde{F}_{h})-F_{\rho}(\theta)||^{2}\big{]}=\operatorname{\mathbb{E}}\big{[}||\tilde{F}_{h}-\operatorname{\mathbb{E}}(\tilde{F}_{h})||^{2}\big{]}+||\operatorname{\mathbb{E}}(\tilde{F}_{h})-F_{\rho}(\theta)||^{2}$ (57)
since the cross term contain $\operatorname{\mathbb{E}}[\tilde{F}_{h}-\operatorname{\mathbb{E}}(\tilde{F}_{h})]$, which equals to zero. We further have
$||\operatorname{\mathbb{E}}(\tilde{F}_{h})-F_{\rho}(\theta)||^{2}\leq\frac{G^{4}\gamma^{2N}}{(1-\gamma)^{2}}\;\text{and}\;\operatorname{\mathbb{E}}\big{[}||\tilde{F}_{h}-\operatorname{\mathbb{E}}(\tilde{F}_{h})||^{2}\big{]}\leq\hat{\sigma}_{F}^{2}$ (58)
@@ Appendix E Proof of Lemma 6
where $(a)$ and $(b)$ follow from $\mu_{F}\leq\left\lVert F_{\rho}(\theta_{k})\right\rVert\leq G^{2}$. Taking conditional expectation on both sides, we obtain
$\operatorname{\mathbb{E}}\left[\left\|\omega_{h+1}-\omega_{k}^{*}\right\|^{2}\right]$ $\leq(1-2\alpha\mu_{F}+2G^{4}\alpha^{2})\|\omega_{h}-\omega_{k}^{*}\|^{2}-2\alpha\langle\omega_{h}-\omega_{k}^{*},\operatorname{\mathbb{E}}\left[\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})\right]\rangle$
$\hskip 28.45274pt+2\alpha^{2}\operatorname{\mathbb{E}}\left\|\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})\right\|^{2}$ (59)
@@
where the last inequality follows from $\left\lVert\omega_{k}^{*}\right\rVert^{2}=\left\lVert F_{\rho}(\theta_{k})^{-1}\nabla_{\theta}J_{\rho}(\theta_{k})\right\rVert^{2}\leq\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}$. Taking expectation yields
$\operatorname{\mathbb{E}}\|\nabla_{\omega}\tilde{L}_{h}$ $-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})\|^{2}$
$\leq 3\operatorname{\mathbb{E}}\|\tilde{F}_{h}-F_{\rho}(\theta_{k})\|^{2}\|\omega_{h}-\omega_{k}^{*}\|^{2}+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\operatorname{\mathbb{E}}\|\tilde{F}_{h}-F_{\rho}(\theta_{k})\|^{2}+3\operatorname{\mathbb{E}}\|\nabla_{\theta}J_{\rho}(\theta_{k})-\tilde{g}_{h}\|^{2}$
$\leq 3\left[\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right]\left\lVert\omega_{h}-\omega_{k}^{*}\right\rVert^{2}+\frac{3\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(3\hat{\sigma}_{g}^{2}+3\hat{\delta}_{g}^{2}\right)$ (60)
@@
$-\langle\omega_{h}-\omega_{k}^{*},$ $\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})]\rangle$
$\leq\frac{\mu_{F}}{4}\|\omega_{h}-\omega_{k}^{*}\|^{2}+\frac{1}{\mu_{F}}\left\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}-F_{\rho}(\theta_{k})(\omega_{h}-\omega_{k}^{*})]\right\|^{2}$
$\leq\frac{\mu_{F}}{4}\|\omega_{h}-\omega_{k}^{*}\|^{2}+\dfrac{1}{\mu_{F}}\left\|\left\{\operatorname{\mathbb{E}}[\tilde{F}_{h}]-F_{\rho}(\theta_{k})\right\}\omega_{h}+\bigg{\{}\nabla_{\theta}J_{\rho}(\theta_{k})-\operatorname{\mathbb{E}}\left[\tilde{g}_{h}\right]\bigg{\}}\right\|^{2}$
@@
where the last inequality follows from $\left\lVert\omega_{k}^{*}\right\rVert^{2}=\left\lVert F_{\rho}(\theta_{k})^{-1}\nabla_{\theta}J_{\rho}(\theta_{k})\right\rVert^{2}\leq\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}$. Substituting the above bounds in (E), we have
$\operatorname{\mathbb{E}}\left[\|\omega_{h+1}-\omega_{k}^{*}\|^{2}\right]$ $\leq\left(1-\frac{3\alpha\mu_{F}}{2}+\dfrac{8\alpha\hat{\delta}_{F}^{2}}{\mu_{F}}+6\alpha^{2}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+2\alpha^{2}G^{4}\right)\|\omega_{h}-\omega_{k}^{*}\|^{2}+\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]$
@@
$\operatorname{\mathbb{E}}[\|\omega_{h+1}-\omega_{k}^{*}\|^{2}]\leq\left(1-{\alpha\mu_{F}}\right)\|\omega_{h}-\omega_{k}^{*}\|^{2}+\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha^{2}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]$
@@
$\operatorname{\mathbb{E}}[\|\omega_{H}-\omega_{k}^{*}\|^{2}]$
$\leq\left(1-{\alpha\mu_{F}}\right)^{H}\operatorname{\mathbb{E}}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\sum_{h=0}^{H-1}\left(1-{\alpha\mu_{F}}\right)^{h}\left\{\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha^{2}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\right\}$
$\leq\exp\left(-{H\alpha\mu_{F}}\right)\operatorname{\mathbb{E}}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\frac{1}{\alpha\mu_{F}}\left\{\dfrac{4\alpha}{\mu_{F}}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha^{2}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\right\}$
$=\exp\left(-{H\alpha\mu_{F}}\right)\operatorname{\mathbb{E}}\|\omega_{0}-\omega_{k}^{*}\|^{2}+\bigg{\{}4\mu_{F}^{-2}\left[2\hat{\delta}_{F}^{2}\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}+\hat{\delta}_{g}^{2}\right]+6\alpha\mu_{F}^{-1}\left[\frac{\mu_{F}^{-2}G^{2}}{(1-\gamma)^{4}}\left(\hat{\sigma}_{F}^{2}+\hat{\delta}_{F}^{2}\right)+\left(\hat{\sigma}_{g}^{2}+\hat{\delta}_{g}^{2}\right)\right]\bigg{\}}$ (62)
@@
$\|\operatorname{\mathbb{E}}[\omega_{h+1}]-\omega_{k}^{*}\|^{2}=\|\operatorname{\mathbb{E}}[\omega_{h}]-\alpha\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-\omega_{k}^{*}\|^{2}$
@@
$=\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*},\operatorname{\mathbb{E}}[{g}_{h}]\rangle+\alpha^{2}\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]\|^{2}$
@@
$\overset{}{=}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*},F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$-2\alpha\langle\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*},\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\rangle+\alpha^{2}\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]\|^{2}$
@@
$\overset{(a)}{\leq}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\mu_{F}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*},\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$\hskip 28.45274pt+2\alpha^{2}\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\|^{2}+2\alpha^{2}\|F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\|^{2}$
@@
$\overset{(b)}{\leq}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\mu_{F}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*},\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$\hskip 28.45274pt+2\alpha^{2}\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\|^{2}+2G^{4}\alpha^{2}\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}$
@@
$\leq(1-2\alpha\mu_{F}+2G^{4}\alpha^{2})\|\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*}\|^{2}-2\alpha\langle\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*},\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\rangle$
@@
$\hskip 28.45274pt+2\alpha^{2}\left\|\operatorname{\mathbb{E}}[\nabla_{\omega}\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\right\|^{2}$ (63)
@@
where $(a)$ and $(b)$ follow from $\mu_{F}\leq\left\lVert F_{\rho}(\theta_{k})\right\rVert\leq G^{2}$. The third term in the last line of (63) can be bounded as follows.
$\|\operatorname{\mathbb{E}}[\nabla_{\omega}$ $\tilde{L}_{h}]-F_{\rho}(\theta_{k})(\operatorname{\mathbb{E}}[\omega_{h}]-\omega_{k}^{*})\|^{2}$
@@
$=\left\|\operatorname{\mathbb{E}}\left[(\tilde{F}_{h}-F_{\rho}(\theta_{k}))(\omega_{h}-\omega_{k}^{*})\right]+(\operatorname{\mathbb{E}}[\tilde{F}_{h}]-F_{\rho}(\theta_{k}))\omega_{k}^{*}+(\nabla_{\theta}J_{\rho}(\theta_{k})-\operatorname{\mathbb{E}}[\tilde{g}_{h}])\right\|^{2}$
@@
$\leq 3\operatorname{\mathbb{E}}\left[\|\operatorname{\mathbb{E}}[\tilde{F}_{h}]-F_{\rho}(\theta_{k})\|^{2}\|\omega_{h}-\omega_{k}^{*}\|^{2}\right]+3\operatorname{\mathbb{E}}\left[\|\operatorname{\mathbb{E}}[\tilde{F}_{h}]-F_{\rho}(\theta_{k})\|^{2}\right]\|\omega_{k}^{*}\|^{2}+3\left\|\nabla_{\theta}J_{\rho}(\theta_{k})-\operatorname{\mathbb{E}}[\tilde{g}_{h}]\right\|^{2}$